Find all x for which x^3-x^2-x+1 /x^3-x^2+x-1 = 0. Please can someone help me? I'm not even sure if this problem belongs in this section, let alone, how to do it. Thanks!!!
Answer by stanbon(75887)

Find all x for which x^3-x^2-x+1 /x^3-x^2+x-1 = 0.
----
Multiply both sides by the denominator to get:
x^3-x^2-x+1 = 0
Factor to get:
(x^2)(x-1) - (x-1) = 0
Factor again:
(x-1)(x^2-1) = 0
Factor again:
(x-1)(x-1)(x+1) = 0
Solutions: x = 1, or x = -1
------------------
But x=1 is extraneous as it makes the original denominator = 0
---------
Only solution: x = -1
==========================
Cheers,
Stan H.